# Dynamic Language Group-Based MoE: Enhancing Efficiency and Flexibility for Code-Switching Speech Recognition

### v1

@@ Abstract
The Mixture of Experts (MoE) approach is ideally suited for tackling multilingual and code-switching (CS) challenges due to its multi-expert architecture. This work introduces the DLG-MoE, which is optimized for bilingual and CS scenarios. Our novel Dynamic Language Group-based MoE layer features a language router with shared weights for explicit language modeling, while independent unsupervised routers within the language group handle attributes beyond language. This structure not only enhances expert extension capabilities but also supports dynamic top-$k$ training, allowing for flexible inference across various top-$k$ values and improving overall performance. The model requires no pre-training and supports streaming recognition, achieving state-of-the-art (SOTA) results with unmatched flexibility compared to other methods. The Code will be released.
@@ 1 Introduction
Thanks to its ability to scale parameters with minimal impact on inference cost, the Mixture of Experts (MoE) approach has achieved exceptional results in natural language processing and computer vision [1, 2, 3]. The inherent multi-expert structure of MoE provides a natural advantage for multilingual tasks. Numerous studies [4, 5, 6] have demonstrated its effectiveness in multilingual automatic speech recognition (ASR), highlighting MoE’s strong performance in this domain. However, these architectures still struggle to effectively handle CS scenarios.
@@
CS scenarios often involve using different languages within utterances, leading to significant challenges in model development due to language confusion. In response, Bi-encoder architectures [7, 8, 9, 10] use separate encoders for each language and employ language-specific masking strategies [8, 9, 11] to enhance individual language processing. However, this approach fails to leverage shared representations across languages, resulting in the loss of cross-lingual contextual information and increased computational complexity. LSR-MoE [12] and LR-MoE [13] mitigate these issues by incorporating two experts within the MoE layer, each dedicated to handling Chinese or English separately, and activate only one expert at a time during training and inference for computational efficiency. Similarly, Google [4] uses multiple experts in the MoE layer, activating two experts to manage multilingual ASR. U2++ MoE [14] highlights MoE’s effectiveness on large-scale Chinese datasets, providing both streaming and non-streaming capabilities within a single MoE and outperforming other models in this aspect.
In this paper, we present DLG-MoE, an efficient and flexible MoE structure. Addressing limitations of previous approaches, we designed language expert groups and a language router with weight sharing at the MoE layer to support flexible expansion. The language router explicitly learns language identification (LID), focusing on distributing frame-level representations to matching language groups based on LID. Within each language group, an unsupervised router specializes in modeling attributes beyond language, such as accents and different domains, distributing inputs to appropriate experts within the group. We also introduce a dynamic top-$k$ routing mechanism during training, allowing for the activation of a variable number of experts during inference. Our contributions are summarized as follows:
  • Our DLG-MoE achieves SOTA performance compared to existing MoE works [7, 12, 13] and also offers streaming recognition capabilities and greater scalability, supporting language-specific and different top-$k$ inference.
@@
  • This paper further investigates the impact of different routing strategies and increasing the number of experts on model performance and also validates the approach on larger datasets and with larger parameters.
Fig. 1: (a) Overall Architecture, (b) DLG-MoE layer, (c) Dynamic Language Group.
@@ 2.1 Overall Model Architecture
The overall architecture of our model is shown in Fig.1(a), comprising Conformer [15] Blocks and DLG-MoE Blocks. The Conformer Block is stacked from vanilla Conformer layers, while the DLG-MoE Block is formed by stacking DLG-MoE layers. Fig.1(b) shows the structure of the DLG-MoE layer, where the second FFN of the Conformer layer is replaced by a dynamic language group. The weight-shared language router (SLR) performs the LID task based on the intermediate representations $h_{inter}$ extracted by the Conformer block and uses the LID results to guide the selection of the appropriate language group.
2.2 Language Router
@@
As illustrated in Fig.1(c), we expand MoE by introducing language expert groups, specifically the Chinese expert group (Zh-Group) and the English expert group (En-Group). Unlike other methods [14, 12, 4], our language router explicitly learns frame-level LID and is trained using an inter-loss function, as expressed by the following formula Equation1.
@@
$\begin{split}\mathcal{L}_{\text{inter}}=&-\left(\log\mathrm{P}_{\mathrm{CTC}}\left(\mathcal{Y}_{\mathrm{LID}}\mid\mathrm{h}_{\text{inter}}\mathrm{W}_{\text{LID}}^{d\times\mathrm{L}+1}\right)\right.\\ &\left.+\log\mathrm{P}_{\mathrm{CTC}}\left(\mathcal{Y}_{\mathrm{ASR}}\mid\mathrm{h}_{\text{inter}}\mathrm{W}_{\text{ASR}}^{d\times V}\right)\right)\end{split}$ (1)
@@
We enhance the SLR’s language routing capability through a multi-task training approach that integrates ASR and LID tasks. Here, $\mathcal{Y}_{\text{ASR}}$ denotes the ground truth for the ASR task, while $\mathcal{Y}_{\text{LID}}$ is derived by mapping the tokens in $\mathcal{Y}_{\text{ASR}}$ to their corresponding language IDs. The linear layers $\mathrm{W}_{\text{LID}}^{d\times\mathrm{L}+1}$ and $\mathrm{W}_{\text{ASR}}^{d\times V}$ represent the transformations, where $d$ is the embedding dimension, $V$ is the vocabulary size, and $L$+1 accounts for the number of languages plus the CTC blank token.
@@
Considering the peaky distribution of CTC which classifies most frames as blank, we use a simple yet empirically effective routing strategy. For the logits from the SLR, which a shape of $\mathbb{R}^{T\times(L+1)}$, where $T$ represents the number of frames, we first remove the blank class to get $\mathbb{R}^{T\times L}$, then select the highest probability among the remaining categories as the language ID for every frame, resulting in a routing table of shape $\mathbb{R}^{T\times 1}$. This greedy strategy, which does not rely on other frames, is effective and beneficial to support streaming ASR. Thus, for the input $h_{pre}$ to the dynamic language group, it is dispatched as $h_{pre}^{zh}$ and $h_{pre}^{en}$ based on the routing information—specifically, the LID results—provided by the SLR, and then processed by the corresponding language expert groups.
2.3 Dynamic Language Group
As shown in Fig. 1(c), once the representations for the same language are obtained, the UnSup-Router within the language expert group focuses on modeling attributes beyond language. For each frame in $h_{pre}^{zh}$, the UnSup-Router selects $k$ experts from the total $n$ experts within the group to collaborate, as described below:
$h_{\text{out }}^{zh}=\sum_{i=1}^{k}G_{i}^{r}E_{i}$ (2)
@@ 2.3 Dynamic Language Group
$G^{r}=\operatorname{Softmax}\left(\operatorname{TopK}\left(h_{pre}^{zh}W_{R}^{d\times n},k\right)\right)$ (3)
@@
Here, $G_{i}^{r}$ and $E_{i}$ denote the representations from the selected experts and their corresponding weights. In Equation 3, $W_{R}^{d\times n}$ is a linear layer that projects the representation $h_{pre}^{zh}$ onto the number of experts $n$ and selects $k$ experts for activation. It is important to note that different frames may be assigned to different sets of $k$ experts, which are then combined in a weighted manner to produce $h_{out}^{zh}$. Subsequently, the Zh-Group and En-Group extract specific language representations $h_{out}^{zh}$ and $h_{out}^{en}$, respectively, combine them to obtain $h_{out}$, and pass it to the next layer. To support various top-$k$ inferences, we employ a dynamic top-$k$ strategy within the expert groups, where the UnSup-Router randomly selects the values of top-k during training. It is crucial to emphasize that neither the language expert groups nor the language router require pre-training; the entire model can be initialized randomly and trained from scratch. All losses are jointly optimized, as illustrated below:
$\mathcal{L}_{\text{total }}=\lambda_{ctc}\mathcal{L}_{\text{ctc }}+(1-\lambda_{ctc})\mathcal{L}_{\text{att}}+\lambda_{\text{inter}}\mathcal{L}_{\text{inter}}$ (4)
@@ 3.1 Experiment setup
Our experiments are conducted on the ASRU-2019 Mandarin-English code-switching challenge dataset [16], along with a subset of 460 hours from Librispeech [17], consistent with [7, 12, 13]. For testing, we use Librispeech-clean (EN), ASRU-Man-test (ZH), and ASRU-CS-test (CS) datasets. Additionally, we have validated our model on an internal multi-accent, multi-domain dataset, which, combined with the ASRU-2019 Challenge dataset, totals 6,000 hours of training data.
@@
We conduct our experiments on the WeNet [18] toolkit. For acoustic features, 80-dimensional log-mel filter banks (FBank) are extracted with a step size of 10ms and a window size of 25ms. All our experimental configurations used a 6-layer Transformer decoder. We set $d^{model}$=256, $n^{heads}$=4 and $d^{ffn}$=2048. Causal convolution was used during training to support streaming. We built Baseline-12 and Baseline-18 with 12 and 18 Conformer layers, respectively. All MoE models had 12 layers, with the first 6 being vanilla Conformer layers and the last 6 incorporating MoE layers. Dense-MoE-4e and Sparse-MoE-4e, which do not include language routing, each had 4 MoE experts. In contrast, DLG-MoE-4e, also with 4 experts, included language routing, allocating two experts per language group. To explore model capacity and different top-$k$ strategies, we created DLG-MoE-8e with 8 experts. For large-scale comparisons, we used $d^{model}$=512 and $n^{heads}$=8 for Baseline-18 and DLG-MoE-8e, while keeping other configurations the same. The hyperparameters $\lambda_{ctc}$ and $\lambda_{inter}$ were set to 0.3 and 0.1, respectively, with the Adam optimizer using $lr$=0.001 and $warmup_{s}teps$=25000. Detailed information can be found in the link provided  [^1] https://github.com/kaihuhuang/Language-Group
3.2 Results of different MoE structures
@@
The results in Table 1 show that MoE structures ($S3$, $S4$) outperform $S2$ with the same number of parameters, indicating that widening the model with MoE might more effective than simply increasing he number of layers. This suggests that even when using only the UnSup-Router, experts can still implicitly learn to model different attributes, such as language or accent, leading to better results. $S4$, which activates only half the experts compared to $S3$, performs better, showing that selective activation of experts is more reasonable. Our proposed $S6$ shows a significant performance improvement over $S3$ and $S4$, proving the effectiveness of language expert groups. The language router’s explicit language differentiation ensures that our language groups focus on specific language representations without interference. In $S7$, without the guidance of an unsupervised router, the expert weights within the language group are assigned the same value, contrasting with the situation in experiment $S6$. In $S6$, the expert weights are dynamically learned through an unsupervised router, highlighting the significant impact of the learned weights in the UnSup-Router. DLG-MoE’s language router focuses on explicit language modeling, allowing the unsupervised routers within the language groups to model attributes beyond language specifically, such as accent or different domains, resulting in better performance. Our model achieves SOTA performance on the CS test set compared to previous MoE works. Specifically, it outperforms Biencoder-MoE(9.76%) [7] , LSR-MoE(10.30%) [12] , and FLR-MoE(9.7%) [13], all using the same training data, while also offering higher flexibility and customizability.
@@
Table 1: Performance of different systems on the test sets. For EN, ZH and CS, we report the word error rate (WER), character error rate (CER), and mixed error rate (MER), respectively. We also assess the model’s computational complexity by evaluating the floating-point operations (FLOPs) for a 20-second input sample.
@@
$S1$ Baseline-12 top1 51M 51M(top-1 decode) 24.8G 6.89 3.35 9.93
$S2$ Baseline-18 top1 72M 72M(top-1 decode) 36.2G 6.52 2.93 9.52
$S3$ Dense-MoE-4e top4 72M 72M(top-2 decode) 34.3G 6.09 2.73 9.36
$S4$ Sparse-MoE-4e top2 72M 60M(top-2 decode) 27.8G 5.99 2.70 9.38
@@
3.3 Analysis of expert capacity and top-k
[…]
Table 2: Results with large-scale datasets and parameters.

### v2

# DYNAMIC LANGUAGE GROUP-BASED MOE: Enhancing Code-Switching Speech Recognition with Hierarchical Routing

@@ Abstract
The Mixture of Experts (MoE) approach is well-suited for multilingual and code-switching (CS) tasks due to its multi-expert architecture. This work introduces the DLG-MoE, a Dynamic Language Group-based MoE optimized for bilingual and CS scenarios. DLG-MoE operates based on a hierarchical routing mechanism. First, the language router explicitly models the language and dispatches the representations to the corresponding language expert groups. Subsequently, the unsupervised router within each language group implicitly models attributes beyond language, and coordinates expert routing and collaboration. The model achieves state-of-the-art (SOTA) performance while also having unparalleled flexibility. It supports different top-$k$ inference and streaming capabilities, and can also prune the model parameters to obtain a monolingual sub-model. The Code will be released.
@@ 1 Introduction
Due to its scalability with minimal impact on inference cost, the Mixture of Experts (MoE) approach has achieved exceptional results in natural language processing and computer vision [1, 2, 3]. The inherent multi-expert structure of MoE provides a natural advantage for multilingual tasks. Numerous studies [4, 5, 6] have demonstrated its effectiveness in multilingual automatic speech recognition (ASR), highlighting MoE’s strong performance in this domain. However, these architectures still struggle to effectively handle CS scenarios.
@@
CS scenarios often involve using different languages within utterances, leading to significant challenges in model development due to language confusion. In response, Bi-encoder architectures [7, 8, 9, 10] use separate encoders for each language and employ language-specific masking strategies [8, 9, 11] to enhance individual language processing. However, this approach fails to leverage shared representations across languages, resulting in the loss of cross-lingual contextual information and increased computational complexity. LSR-MoE [12] and LR-MoE [13] mitigate these issues by incorporating two experts within the MoE layer, each dedicated to handling Chinese or English separately, and activate only one expert at a time during training and inference for computational efficiency. Similarly, Google [4] uses multiple experts in the MoE layer, activating two experts to manage multilingual ASR. In addition, there are some other MoE methods have also demonstrated the potential of MoE for Mandarin [14, 15] as well as multi-accent and multi-domain scenarios [16].
Fig. 1: (a) Overall Architecture, (b) DLG-MoE layer, (c) Dynamic Language Groups.
The proposed DLG-MoE in this work aims to address the limitations of existing MoE methods in handling CS tasks, while fully leveraging the advantage of MoE’s scalability in parameters. Our hierarchical routing mechanism primarily relies on the language router and the unsupervised router within each language group. The language router focuses on language-based routing and distributing features to language group to capture language-specific representations. While the unsupervised router within the language group focuses on implicitly modeling accents and domains beyond language. We also introduce a dynamic top-$k$ training strategy, allowing for the activation of a variable number of experts during inference. Our contributions are summarized as follows:
  • The proposed hierarchical routing mechanism and dynamic language groups in DLG-MoE achieves state-of-the-art (SOTA) performance on CS tasks, surpassing existing related MoE methods[7, 12, 13].
  • The proposed approach is more scalable in terms of expert capacity, and also supports flexible top-$k$ inference as well as streaming recognition capabilities.
@@
  • This work thoroughly examines the impact of different routing strategies and the number of experts on model performance. Furthermore, we have validated the effectiveness of our methods on larger datasets and with more model parameters.
@@ 2.1 Overall Model Architecture
The overall architecture of our model is shown in Fig.1(a), comprising Conformer [17] Blocks and DLG-MoE Blocks. The Conformer Block is stacked from vanilla Conformer layers, while the DLG-MoE Block is formed by stacking DLG-MoE layers. Fig.1(b) shows the structure of the DLG-MoE layer, where the second FFN of the Conformer layer is replaced by dynamic language groups. The weight-shared language router (SLR) performs the frame-wise LID task based on the intermediate representations $\bm{H}_{\text{inter}}$ extracted by the Conformer block and uses the LID results to guide the selection of the appropriate language group.
2.2 Shared Language Router
@@
As shown in Fig. 1(c), the weight-shared language router first routes the features to the corresponding language group based on their language attributes. Unlike other methods [14, 12, 4], our language router explicitly learns frame-level LID and is trained using an inter-loss function, as expressed by the following formula Equation1.
@@
$\begin{split}\mathcal{L}_{\text{inter}}=&-\left(\log\mathrm{P}_{\mathrm{CTC}}\left(\mathcal{Y}_{\mathrm{LID}}\mid\bm{H}_{\text{inter}}\mathbf{W}_{\text{LID}}^{D\times\mathrm{L}+1}\right)\right.\\ &\left.+\log\mathrm{P}_{\mathrm{CTC}}\left(\mathcal{Y}_{\mathrm{ASR}}\mid\bm{H}_{\text{inter}}\mathbf{W}_{\text{ASR}}^{D\times V}\right)\right)\end{split}$ (1)
@@
We enhance the SLR’s language routing capability through a multi-task training approach that integrates ASR and frame-level LID tasks. Here, $\mathcal{Y}_{\text{ASR}}$ denotes the ground truth for the ASR task, while $\mathcal{Y}_{\text{LID}}$ is derived by mapping the tokens in $\mathcal{Y}_{\text{ASR}}$ to their corresponding language IDs. The $\mathbf{W}_{\text{LID}}^{D\times\mathrm{L}+1}$ and $\mathbf{W}_{\text{ASR}}^{D\times V}$ represent the weights of the linear layer, where $D$ is the embedding dimension, $V$ is the vocabulary size, and $L$+1 accounts for the number of languages plus the $blank$ token. The key point to emphasize is that frame-level LID labels are not necessary, as the Connectionist Temporal Classification (CTC) can align the frames to the LID label sequence $\mathcal{Y}_{\text{LID}}$ automatically.
@@
Given the peaky distribution of CTC, which classifies most time steps as $blank$, we use a simple yet effective routing strategy. Specifically, for each time step, we select the index with the maximum value among the $non$-$blank$ language categories as the language ID. This greedy strategy, which does not rely on other time steps, is effective and beneficial to support streaming ASR. The input $\bm{H_{\text{pre}}}$ is then dispatched to the corresponding language expert groups ($\bm{H_{\text{pre}}^{\text{ZH}}}$ and $\bm{H_{\text{pre}}^{\text{EN}}}$) based on the language ID provided by the SLR.
2.3 Dynamic Language Groups
Then within each language group, the UnSup-Router performs second-level routing in an unsupervised manner. At this stage, the routing is no longer based on language, but instead implicitly models and routes based on other attributes beyond just language. For each time step in $\bm{H_{\text{pre}}^{\text{ZH}}}$, the UnSup-Router will selects $k$ experts from the total $n$ experts within the group to collaborate, as described below:
$\bm{H_{\text{out}}^{\text{ZH}}}=\sum_{i=1}^{k}\mathbf{G}_{i}^{r}\bm{E_{i}}$ (2)
@@ 2.3 Dynamic Language Groups
$\mathbf{G}_{i}^{r}=\operatorname{Softmax}\left(\operatorname{TopK}\left(\bm{H_{\text{pre}}^{\text{zh}}}\mathbf{W}_{R}^{D\times n},k\right)\right)$ (3)
@@
Here, $E_{i}$ and $\mathbf{G}_{i}^{r}$ denote the outputs from the selected experts and their corresponding weights. In Equation 3, $\mathbf{W}_{R}^{D\times n}$ is a linear layer that projects the representation $\bm{H}_{\text{pre}}^{\text{ZH}}$ onto the number of experts $n$ and selects $k$ experts for activation. It is important to note that different time steps may be assigned to different sets of $k$ experts, which are then combined in a weighted manner to produce $\bm{H}_{\text{out}}^{\text{ZH}}$. Subsequently, the ZH-Group and EN-Group extract language-specific representations $\bm{H}_{\text{out}}^{\text{ZH}}$ and $\bm{H}_{\text{out}}^{\text{EN}}$, respectively, combine them to obtain final output $\bm{H}_{\text{out}}$, and pass it to the next layer. To support various top-$k$ inferences, we employ a dynamic top-$k$ strategy within the language groups, where the UnSup-Router randomly selects the values of top-$k$ during training. It is crucial to emphasize that neither the language groups nor the language router require pre-training; the entire model can be initialized randomly and trained from scratch. All losses are jointly optimized, as illustrated below:
$\mathcal{L}_{\text{total }}=\lambda_{\text{ctc}}\mathcal{L}_{\text{ctc }}+(1-\lambda_{\text{ctc}})\mathcal{L}_{\text{att}}+\lambda_{\text{inter}}\mathcal{L}_{\text{inter}}$ (4)
@@ 3.1 Experiment Setup
Our experiments are conducted on the ASRU-2019 Mandarin-English code-switching challenge dataset [18], along with a subset of 460 hours from Librispeech [19]. The total duration of the datasets used in our experiments is approximately 1100 hours, consistent with previous works [7, 12, 13]. For testing, we use Librispeech-clean (EN), ASRU-Man-test (ZH), and ASRU-CS-test (CS) datasets. Additionally, we have validated our model on an internal multi-accent, multi-domain dataset, which, combined with the ASRU-2019 Challenge dataset, totals 6,000 hours of training data.
@@
We conduct our experiments on the WeNet [20] toolkit. For acoustic features, 80-dimensional log-mel filter banks (FBank) are extracted with a step size of 10ms and a window size of 25ms. All our experimental configurations used a 6-layer Transformer decoder. We set $d^{model}$=256, $n^{heads}$=4 and $d^{ffn}$=2048. Causal convolution was used during training to support streaming. We built Baseline-12 and Baseline-18 with 12 and 18 Conformer layers, respectively. All MoE models had 12 layers, with the first 6 being vanilla Conformer layers and the last 6 incorporating MoE layers. Dense-MoE-4e and Sparse-MoE-4e, which do not include language routing, each had 4 experts. In contrast, DLG-MoE-4e, also with 4 experts, included language routing, allocating two experts per language group. To explore model capacity and different top-$k$ strategies, we created DLG-MoE-8e with 8 experts. For large-scale comparisons, we used $d^{model}$=512 and $n^{heads}$=8 for Baseline-18 and DLG-MoE-8e, while keeping other configurations the same. The hyperparameters $\lambda_{ctc}$ and $\lambda_{inter}$ were set to 0.3 and 0.1, respectively, with the Adam optimizer using $lr$=0.001 and $warmup\_steps$=25000. Detailed information can be found in the link provided [^1] https://github.com/kaihuhuang/Language-Group.
3.2 Analysis of Results for Different MoE Structures
@@
The results in Table 1 show that MoE structures ($S3$, $S4$) outperform $S2$ with the same parameter count, suggesting that widening the model with MoE may be more effective than simply increasing the number of layers. This implies that even using only the UnSup-Router, experts can still implicitly learn to model different attributes, such as language or accent, resulting in better performance. $S4$, which activates only half the experts compared to $S3$, performs better, indicating that selective expert activation is more reasonable.
Our proposed $S6$ shows a significant performance improvement over $S3$ and $S4$, proving the effectiveness of language expert groups. The language router’s explicit language differentiation ensures that language groups focus on specific language representations without interference by other language domains. In $S7$, without the unsupervised router, the expert weights within the language group are equally assigned, contrasting $S6$’s dynamically learned expert weights through the UnSup-Router (9.23$\rightarrow$9.09%), highlighting the effect of UnSup-Router. The comparative results of $S4$, $S6$, and $S7$ demonstrate the hierarchical routing approach’s effectiveness in DLG-MoE. The language router and UnSup-Router have accurately modeled language and other attributes, respectively. Our model achieves SOTA compared to previous MoE works in CS tasks, outperforming Biencoder-MoE (9.76%)[7], LSR-MoE (10.30%)[12], and FLR-MoE (9.7%)[13], using the same training data, while offering higher flexibility and customizability.
@@
Table 1: The performance of various systems on English (EN), Chinese (ZH), and Code-switching (CS) test sets is evaluated. We report Word Error Rate (WER), Character Error Rate (CER), and Mixed Error Rate (MER) for each scenario. Additionally, computational complexity is measured by the number of Floating-Point Operations (FLOPs) required for a 20-second input.
@@
$S1$ Baseline-12 top-1 51M 51M(top-1 decode) 24.8G 6.89 3.35 9.93
$S2$ Baseline-18 top-1 72M 72M(top-1 decode) 36.2G 6.52 2.93 9.52
$S3$ Dense-MoE-4e top-4 72M 72M(top-2 decode) 34.3G 6.09 2.73 9.36
$S4$ Sparse-MoE-4e top-2 72M 60M(top-2 decode) 27.8G 5.99 2.70 9.38
@@
3.3 Analysis of Expert Capacities and Top-k
In this section, we analyze the impact of different top-$k$ training strategies and increasing the number of experts on the performance of DLG-MoE. Contrary to the conclusions from LSR-MoE [12], we find that increasing the number of experts can yield better results (9.10$\rightarrow$8.92%). This discrepancy is likely because LSR-MoE did not incorporate language routing, making implicit language modeling more challenging with additional experts.
The results in $S9$-$S11$ demonstrate that using top-2 training and inference generally yields better performance, while inconsistent training and inference top-$k$ values lead to a significant performance drop. This issue is resolved by employing dynamic top-$k$ training. Surprisingly, the dynamic top-$k$ model $S11$ even outperformed the fixed top-$k$ trained models $S9$ and $S10$, likely due to the dynamic top-$k$ approach enhancing the model’s generalization. However, the dynamic top-$k$ model $S5$ performed worse than $S4$, likely because the dynamic top-$k$ approach exacerbated the difficulty of implicit language modeling in the router in $S5$, while the hierarchical routing mechanism in DLG-MoE avoids this issue. Our system $S11$ maintains the same number of activated parameters and similar FLOPs as $S1$ in top-1 decoding, yet achieves up to 22.2%, 26.9%, and 8.9% relative improvement in EN, ZH, and CS, respectively. Our method continues to show significant advantages on larger datasets and with more parameters, as shown in Table 2.
Table 2: Results with larger datasets and parameters.